The lengths of the sides of a triangle are three consecutive whole numbers. The perimieter is 153cm. Determine the length of the three sides and give a complete algebraic solution? Thanks everyone!!
Answer by rfer(16322)

x+x+1+x+2=153
3x+3=153
3x=150
x=50
x+1=51
x+2=52